A steam boiler is to have a 300-mm-diameter hand hole placed on one end. If the maximum internal pressure is to be 1,3 MPa, how many 15-mm bolts will be required to keep the cover plate in position? The root diameter of a 15-mm bolt is 12 mm. The allowable tensile stress in the bolts is 100 MPa and the initial stress in the bolts when they are tightened is 60 MPa.
1
Expert's answer
2020-12-07T03:48:56-0500

Tightening stress + Additional stress on bolt = Allowable stress for bolt

60+ "\frac{F}{Ab} = 100" ( here Ab = root area of bolt)

60 + "\frac{F}{(\pi)/{4}\times 12^2}" = 100


solving we get

F = 4523.89342 N

Now

F' = P (A) [ pressure p active on area A]

F' = 1.3"\times \frac{\pi}{4}" "\times 300^2"

F' = 91891.58512 N

n = "\frac{F'}{F}"


n= "\frac{91891.58512}{4523.89342}"


n= 20.31250

so in whole number n=21 bolts are required